OpenStudy (anonymous):

sin theta = -4/5 cos beta = -5/13 I have to find cos theta/2 I don't think my final answer is correct so may someone go through it with me

zepdrix (zepdrix):

So what did you get for your cos(theta) so far? :o

zepdrix (zepdrix):

This problem is confusing, I'm not even sure how the cos(beta) is going to help us...

OpenStudy (anonymous):

3/5

zepdrix (zepdrix):

Ok good. And then we use our Half-Angle Identity or something, ya?\[\Large\rm \cos\frac{\theta}{2}=\sqrt{\frac{1+\cos \theta}{2}}\]

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

\[\Large\rm \cos\frac{\theta}{2}=\sqrt{\frac{1+\frac{3}{5}}{2}}\]

OpenStudy (anonymous):

then I multiplied 5 to he 1 and 2 and left 3

zepdrix (zepdrix):

Multiplied by 5? Like top and bottom by 5? It should give you something like this if you do that:\[\Large\rm \cos\frac{\theta}{2}=\sqrt{\left(\frac{1+\frac{3}{5}}{2}\right)\frac{5}{5}}\]gives,\[\Large\rm =\sqrt{\frac{5+3}{2\cdot5}}\]

OpenStudy (anonymous):

yep

zepdrix (zepdrix):

I guess they'll probably want you to rationalize your answer, so let's do that, it'll look a little nicer getting the square root out of the bottom.

zepdrix (zepdrix):

\[\Large\rm =\sqrt{\frac{8}{10}}=\sqrt{\frac{4}{5}}=\frac{\sqrt4}{\sqrt5}\]

zepdrix (zepdrix):

We'll multiply top and bottom by sqrt(5), ya?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

sqrt of 4 is 2

zepdrix (zepdrix):

\[\Large\rm =\frac{2}{\sqrt{5}}\cdot\frac{\sqrt5}{\sqrt5}=?\]

OpenStudy (anonymous):

\[2\sqrt{5}/5\]

zepdrix (zepdrix):

Yayyy good job \c:/

OpenStudy (anonymous):

Thank you again haha

zepdrix (zepdrix):

Maybe we want a plus/minus on the root, they don't really give us enough information to determine if it should be positive or negative.. hmm

OpenStudy (anonymous):

I'm going over a review so I'm trying to make sure everything is right

zepdrix (zepdrix):

ah neato ^^